# EAGLE: The First Event Camera Dataset Gathered by an Agile Quadruped Robot

### v1

@@ IV-B Indoor Sequences
The indoor sequences were collected in 13 diverse environments. In the dining hall, building floor, and home environments, we recorded two sequences per site, with each sequence showcasing a different set of quadruped gaits. For other indoor environments where we can control the lighting condition, data collection was expanded to encompass a range of lighting conditions – well-lit, dark, HDR, and blinking – while consistently applying two sets of quadruped gaits within each lighting condition. Specifically, in laboratory environments, four sequences were recorded in one site under both consistently well-lit and HDR lighting conditions. To make HDR condition, we used bright light sources in a dark environment. In classroom and kitchen environments, we gathered six sequences in each place across well-lit, HDR, and blinking lighting conditions. In the MoCap space, we arranged various objects on the ground to create three different scenes, recording six sequences in each environment under well-lit, dark, and blinking lighting conditions.
Indoor sequences, similar to outdoor ones, also encounter challenges such as diverse lighting conditions, dynamic objects, and slippery floors. Additionally, indoor sequences introduce unique challenges: light reflections from surfaces such as floors and glasses can complicate perception, and closely positioned obstacles may briefly obstruct cameras’ view, adding extra complexity to navigation and sensing.
| Faster-LIO | VINS-Fusion | USLAM | E-ATS | |||||||||
| mocap_env1_trot (Fig. 6(a)) | 4.43 | 0.59 | 1.15 | |||||||||
| mocap_env1_comb (Fig. 6(b)) | 10.72 | 0.60 | 1.41 | |||||||||
| mocap_env1_dark_trot (Fig. 6(c)) | 0.86 | 2.78 | 3.60 | |||||||||
| mocap_env1_dark_comb | 3.35 | 5.91 | 0.75 | |||||||||
| mocap_env1_blink_trot (Fig. 6(d)) | 0.99 | 3.24 | 5.52 | |||||||||
| mocap_env1_blink_comb | 3.12 | 6.36 | 1.03 | |||||||||
| env5_backflip1 (Fig. 7) | ||||||||||||
| lab1_hdr_comb (Fig. 6(e)) | 1.10 | 6.35 | 250.26 | |||||||||
| learning_center_trot (Fig. 6(f)) | 0.75 | 5.58 | 104.09 | |||||||||
| sidewalk1_day_trot (Fig. 6(g)) | 14.40 | 1.00 | ||||||||||
| downtown1_night_comb (Fig. 6(h)) | 0.77 | 7.86 | 224.93 | |||||||||
| residential_area_night_comb | ||||||||||||
The symbol / differentiates evaluations on partial and full trajectories, with partial ones manually truncated and evaluated before significant divergence. The symbol means not included. RPE for backflip sequences is assessed at 0.1-second intervals.
@@ IV-C Backflip Sequences
The backflip, demonstrating the robot’s dynamic capabilities with rotations up to , was captured in 7 indoor environments and 1 outdoor environment, with two sequences in each environment. We excluded the LiDAR sensor due to its substantial weight, which can bother the success of the acrobatic movement, and the significant motion distortion in the point cloud data. Each sequence contains three consecutive backflips, after which the robot returns to the initial marked position.
Backflip sequences present distinct challenges that are different from indoor and outdoor sequences. The backflip motion’s high speed lead to significantly blurred images from the RealSense RGB camera. Ground contact causes abrupt changes in acceleration and angular velocity (Fig. 1 and Fig. 4), posing significant challenges for visual-inertial systems and serving as a crucial test for their adaptability in highly dynamic robot motions. Another unique challenge in backflip sequence is the difficulty in feature detection. Although we chose environments that offer visual features from multiple angles, captured images may appear featureless when the camera’s view is directed toward the bare ceiling or floor.
IV-D Limitations
During data gathering, we observed two artifacts that could be potential limitations of the dataset. The first is occasional missing event data in outdoor daytime sequences, particularly in environments with complex textures, attributed to the limited throughput of the DAVIS346 event camera (RGB + Event). The second artifact is blinking dot noise observed in the dark regions of RGB images captured by the RealSense camera. This noise becomes apparent when the camera’s auto-exposure feature is enabled to optimize performance in high dynamic range environments.
@@ V Evaluation
To assess the difficulty of the dataset and the quality of the parameters, we tested various sequences with state-of-the-art algorithms for several sensor pairs, including Faster-LIO (LiDAR + IMU) [29], E-ATS (Event + RGB-D) [30], UltimateSLAM (Event + RGB + IMU, noted as USLAM) [19], and VINS-Fusion (RGB + IMU) [31]. USLAM operates in event-IMU mode for indoor evaluations due to the lack of RGB images from the event camera, and switches to event-RGB-IMU mode for outdoor environments.
For sequences captured within the MoCap environment (labeled as mocap_*), we use ground-truth poses from the MoCap system for evaluation. For sequences without a MoCap system, we use the pose estimated by the Faster-LIO algorithm as ground truth because the estimation results of the Faster-LIO algorithm demonstrate good matches between the initial and final poses, as shown in Fig. 6. Also, the Faster-LIO shows superior accuracy on indoor MoCap sequences as evidenced in Table III. We evaluate each algorithm based on two key metrics: Absolute Pose Error (APE) and Relative Pose Error (RPE). For RPE computation, rotation is measured in degrees per second and translation is measured in centimeters per second, except for backflip sequences where we measure in degrees and centimeters per 0.1 seconds because the duration of one backflip motion is short. APE is measured in centimeters to evaluate pose estimation accuracy. The evaluation is based on the open-source software EVO333https://github.com/MichaelGrupp/evo. The evaluation results are summarized in Table III to provide a quantitative comparison.
Fig. 6 and Fig. 7 offer a qualitative visualization. The results show that most algorithms achieve accurate pose estimation with trot gait and a well-lit lighting condition (Fig. 6 (a)). However, pose estimation accuracy drops with more dynamic gaits such as bounding and pronking gaits (Fig. 6 (b)), and in challenging dark or blinking lighting conditions (Fig. 6 (c, d)). Notably, backflip sequences, containing three consecutive backflips, present significant challenges that lead to the divergence of all tested algorithms. Consequently, our evaluation focuses on the accuracy within a single backflip to assess algorithm performance during the extreme dynamic motion (Fig. 7). Among the three tested algorithms, only one (E-ATS) survived from divergence. Our evaluation results imply that the dataset includes sufficiently challenging scenarios for state-of-the-art visual-inertial algorithms and can serve as valuable assets for future benchmarks of advanced event-based visual-inertial odometry algorithms.
@@ VI CONCLUSIONS
In this paper, we introduce the EAGLE dataset, the pioneering dataset designed to investigate the utilization of an event camera, RGB-D camera, LiDAR, IMU, and joint encoders on an agile quadruped robot. The EAGLE dataset offers a rich collection of indoor and outdoor sequences under various quadruped gaits and lighting conditions. Additionally, it includes backflip sequences that highlight the agility of the robot. We envision the EAGLE dataset as a cornerstone for propelling forward research in event camera applications in the field of legged robotics.
@@ ACKNOWLEDGMENT
We express our gratitude to Naver Labs and MIT Biomimetic Robotics Lab for providing the Mini-cheetah robot as a research platform for conducting dynamic motion studies on legged robots.

### v3

@@ IV-B Indoor Sequences
[…]
In addition to the challenges similar to outdoor ones such as diverse lighting conditions, dynamic objects, and slippery floors, indoor environments introduce unique complications such as the reflections from floors or glasses that can confuse perception, and closely positioned obstacles that may briefly obstruct cameras’ view.
@@ IV-C Backflip Sequences
The backflip, involving the rotations up to , was captured in 7 indoor and 1 outdoor environments, with two sequences in each environment. We excluded the LiDAR sensor due to its substantial weight, which can bother the success of the acrobatic movement, and the significant motion distortion in the point cloud data. Each sequence contains three consecutive backflips and the robot returns to the initial starting position after the three flips.
Backflip sequences present distinct challenges that are different from other sequences. The backflip motion’s high speed leads to significantly blurred images from the RealSense RGB camera. Ground contact at landing causes abrupt changes in acceleration and angular velocity (Fig. 1 and Fig. 4), posing significant challenges for visual-inertial systems and serving as a crucial test for their robustness in agile robot motions. In addition, feature detection becomes also difficult during the acrobatic motion. Although we chose environments that offer visual features from multiple angles, captured images may appear featureless when the camera’s view is directed toward the bare ceiling or floor.
Faster-LIO | VINS-Fusion | ORB-SLAM3 | USLAM | E-ATS | ||||||||||
Seq.1 Fig.6a | 0.6/2.5 | |||||||||||||
Seq.2 Fig.6b | / | / | ||||||||||||
Seq.3 Fig.6c | 1.3/7.4 | |||||||||||||
Seq.4 | 2.1/8.6 | |||||||||||||
Seq.5 Fig.6d | 2.8/6.0 | |||||||||||||
Seq.6 | 1.4/8.6 | |||||||||||||
Seq.7 Fig.7 | ||||||||||||||
Seq.8 Fig.6e | ||||||||||||||
Seq.9 Fig.6f | ||||||||||||||
Seq.10 Fig.6g | ||||||||||||||
Seq.11 Fig.6h | 0.9/9.5 | 6.1/476.2 | ||||||||||||
Seq.12 |
From top to bottom, the sequences are as follows: Seq. 1-6 (mocap environment) depicts various gaits under different lighting conditions: trot gait/well-lit, combined gaits/well-lit, trot gait/dark, combined gaits/dark, trot gait/blinking light, and combined gaits/blinking light. Seq. 7 shows a backflip and Seq. 8 is a trot gait in a laboratory with HDR lighting. Seq. 9 is a trot gait in a well-lit learning center. Seq. 10 shows a trot gait on a sidewalk during the daytime. Seq. 11 features combined gaits at night in a downtown setting while Seq. 12 shows combined gaits at night in a residential area. The symbol * indicates a significant divergence, which means that the trajectory diverges by 10% of the total trajectory length. In this case, we truncated the trajectory and evaluate only the portion of the data before the divergence. The symbol means not included. The term is used to denote ground truth. RPE for backflip sequences (Seq.7) is assessed at 0.1-second intervals. USLAM operates in three modes: event + constructed frames + IMU ( ), event + frame + IMU ( ), or event + IMU (without and ).
@@ V Evaluation
To assess the difficulty and quality of the dataset, we tested various sequences with state-of-the-art algorithms for several sensor pairs, including Faster-LIO (LiDAR + IMU) [31], E-ATS (Event + RGB-D) [32], UltimateSLAM (Event + RGB + IMU, noted as USLAM) [21], VINS-Fusion (RGB + IMU) [33], ORB-SLAM3 (RGB-D with loop closure) [34]. All algorithms are tested using open-source implementation provided by the authors, but with CEAR’s intrinsic/extrinsic parameters. We diligently tuned the parameters of each algorithm to maximize the performance and found the different sets for indoor and outdoor environments. In the case of USLAM algorithm, we utilize three different setups. When using DVXplorer Lite, which lacks RGB images, we construct grayscale images from event data by utilizing E2VID [22] (event + constructed frame + IMU mode). When RGB image is available (DAVIS346), we use default setting (event + frame + IMU mode). When the environment is dark, we use only event and IMU since the quality of constructed images was low (event + IMU mode). In the evaluation, we use the best results among the ones from three modes.
[…]
The results show that most algorithms achieve accurate pose estimation with trot gait in well-lit conditions. However, non-event-based algorithms’ accuracy drops with dynamic gaits like bounding and pronking, while event-based methods remain stable (Table III Seq.1 and Seq.2). Challenging lighting conditions (dark and blinking) further affect accuracy (Table III Seq.3-6). E-ATS performs poorly in Seq.3-6 due to darkness failing to trigger event data and blinking light breaking the intensity-constant assumption. Notably, backflip sequences, containing three consecutive backflips, present significant challenges, so none of the tested algorithms can handle more than one backflip. Consequently, we limit our evaluation to a single backflip(Fig. 7). Among the four tested algorithms, only one (E-ATS) survived from divergence. Our evaluation results imply that the dataset includes sufficiently challenging scenarios for state-of-the-art algorithms and serve as valuable assets for future benchmarks of advanced event-based visual-inertial odometry algorithms.
@@ VI CONCLUSIONS
In this paper, we introduce the CEAR, the pioneering dataset designed to investigate the perception system of a dynamic quadruped robot by utilizing event/RGB-D cameras, LiDAR, IMU, and joint encoders. The CEAR dataset offers a comprehensive collection of indoor and outdoor sequences under various quadruped gaits and lighting conditions. Additionally, it includes backflip sequences to reflect the unique challenges in motion estimation of dynamic legged robots. We envision the CEAR dataset as a cornerstone for exploring the rapid perception of event cameras in robotics research. Furthermore, we anticipate that the dataset will enable the exploitation of unique dynamics introduced by quadruped robots and facilitate the development of specialized algorithms for legged systems. In future work, we plan to add a hardware synchronization module and collect data under more versatile, dynamic movements and develop data-driven methods using CEAR dataset.
@@ ACKNOWLEDGMENT
We express our gratitude to Naver Labs and MIT Biomimetic Robotics Lab for providing the Mini-cheetah robot as a research platform.